OpenStudy (anonymous):

A simple pendulum suspended to the ceiling of a moving car on a horizontal road is aligned at 27 degree with the vertical. If it started from rest , what might be the distance travelled by it as it acquires a velocity of 36 kmph? ( Tan 27 = 0.5 and g =10m/s) options-10m, 15,5 &20m.

OpenStudy (anonymous):

 now use Vf^2 = Vi^2 +2ad for Vf, convert 36 kmph to m/s Vi = 0 a= .5g find d

OpenStudy (anonymous):

(36x1000/3600)10 m/s^2 =2 x0.5g xd or d=100/10 or d= 10 m. Thanks